# Neutrino reheating predictions with non-thermal leptogenesis

### v1

@@ 4.4 Example 4: Natural inflation
$V=\Lambda^{4}\left(1+\cos\phi/f\right)\,.$ (46)
@@
Planck 2018 central value of $n_{s}=0.9649$ gives $f>5.34M_{\mathrm{P}}$. We can express $H_{k}$ as a function of $n_{s}$ with the help of eq. (52) and eq. (54). Similarly, we can get $V_{\mathrm{end}}$ as a function of $n_{s}$ with eq. (50) eq. (51) and $H_{k}$. Note that all these quantities are also functions of $f$, for which we choose two benchmark values, $f=6M_{\mathrm{P}}$ and $f=10M_{\mathrm{P}}$. The overall scale $\Lambda$ can be fixed by the scalar perturbation amplitude, which is at $\mathcal{O}(10^{16})$ GeV. The inflaton mass is $M_{\phi}=\Lambda^{2}/f$, which is $1.34\times 10^{13}$ GeV for $f=6M_{\mathrm{P}}$ ($1.47\times 10^{13}$ GeV for $f=10M_{\mathrm{P}}$).
@@
Following the $N_{\mathrm{RH}}$-$N_{k}$ relation in (10), we find a prediction of $n_{s}=0.958$ for $f=6M_{\mathrm{P}}$ and $n_{s}=0.964$ for $f=10M_{\mathrm{P}}$; the latter is within the current $1\sigma$ region. $N_{\mathrm{RH}}=0$ also allows us to equate $\rho_{\mathrm{end}}$ with $\rho_{\mathrm{RH}}$. Hence, we find a relation of $T_{\mathrm{RH}}$ with $n_{s}$. The predicted $n_{s}$ corresponds to $T_{\mathrm{RH}}=2.63\times 10^{15}$ GeV for $f=6M_{\mathrm{P}}$ and $T_{\mathrm{RH}}=2.76\times 10^{15}$ GeV for $f=10M_{\mathrm{P}}$. The generated baryon asymmetry are $1.04\times 10^{-4}$ ($f=6M_{\mathrm{P}}$) and $1.00\times 10^{-4}$ ($f=10M_{\mathrm{P}}$), both are several orders larger than needed.
@@
In this scenario, $N_{\mathrm{RH}}$ calculated from (18) and (19) depends on five parameters:
$M_{\phi},~{}M_{N},~{}T_{\mathrm{RH}},~{}n_{s},~{}f\,$
We also choose to work with $f=6M_{\mathrm{P}},10M_{\mathrm{P}}$ and $M_{N}=10^{11},10^{9}$ GeV. Then the $N_{\mathrm{RH}}$-$N_{k}$ relation allows us to determine $T_{\mathrm{RH}}$ versus $n_{s}$. From (21) we can find an upper limit for $n_{s}$. Following the arguments in [11], we find a lower limit for $n_{s}$ from the requirement $N_{k}\geq 29$. Note that for different $f$, the upper/lower limits for $n_{s}$ are different.
@@
Figure 7: Predictions in natural inflation with neutrino reheating in $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. Other shadings are the same as in the figure 1.
Figure 8: Predictions in natural inflation with neutrino reheating in $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. The solid lines are predictions on the tensor-to-scalar ratio, while the dotted lines are predictions on the baryon asymmetry. The blue lines correspond to $f=6\mathrm{M}_{\mathrm{P}}$ and the red lines are $f=10\mathrm{M}_{\mathrm{P}}$. Other shadings are the same as in the figure 2.
@@
In figure 7, we present various quantities as a function of the spectral index. We see that lines with $f=10M_{\mathrm{P}}$ overlap with the observational region, while lines with $f=6M_{\mathrm{P}}$ do not overlap at all. A larger $f$ corresponds to a lower $T_{\mathrm{RH}}$, a smaller $N_{k}$ and a smaller $\omega_{\mathrm{RH}}$. Different $M_{N}$ lines converge in the $\omega_{\mathrm{RH}}$ $n_{s}$ plane. As the Universe approaches radiation domination ($\omega=1/3$), the role of $M_{N}$ becomes increasingly irrelevant. Our results confirm that requiring $\omega_{\mathrm{RH}}\leq 1/3$ gives $r\geq 0.05$ [11], and are compatible with updated results in [42].
In figure 8, we show the predictions of $Y_{\mathrm{B}}$ in the $n_{s}$ $r$ plane. Since $Y_{\mathrm{B}}$ is degenerate in $M_{N}$ and $r$ is independent of $M_{N}$, there are only two lines (corresponding to $f=6M_{\mathrm{P}},10M_{\mathrm{P}}$) for each quantity. The solid lines are predictions on $r$. The $r$ line corresponding to $f=6M_{\mathrm{P}}$ overlaps with the combined Planck 2018 $+$ BICEP/Keck $2\sigma$ region, while $f=10M_{\mathrm{P}}$ does not. As for the predictions on $Y_{\mathrm{B}}$, both lines intersect the observation band within $n_{s}$ $1\sigma$ region. Although we take only two benchmark values of $f$, it is conceivable that other values of $f$ could predict $r$ within the allowed observation region. This degeneracy can be alleviated when $Y_{\mathrm{B}}$ is considered. As the baryon asymmetry changes rapidly with $n_{s}$, it will help to break the degeneracy by giving very distinctive values of $n_{s}$. Large $f$ predicts $r$ value outside the $2\sigma$ observation region. Taking into account other effects such as dissipative effects [51, 52] or modified gravity  [53, 54, 55, 56, 57, 58] can suppress $r$ and render a natural inflation model compatible with all the constraints.
@@ 5 Discussions and conclusions
In this work, we continue investigating the intriguing possibility of connecting inflation and neutrino physics through non-thermal leptogenesis [1]. We assume a direct inflaton-RHN coupling. The RHNs decay reheats the Universe through their subsequent decays to SM particles and we called it “neutrino reheating”. Depending on the relative sizes of $\Gamma_{N}$ and $\Gamma_{\phi}$, we consider two scenarios of neutrino reheating. In $\Gamma_{N}\gg\Gamma_{\phi}$ scenario, the RHNs decay instantly and correspond to $N_{\mathrm{RH}}=0$. We find a prediction on $n_{s}$, $T_{\mathrm{RH}}$ and hence $Y_{\mathrm{B}}$ for each model. The predicted $Y_{\mathrm{B}}$ are at $\mathcal{O}(10^{-4})$ to $\mathcal{O}(10^{-5})$, which are much larger than the observed value and leave much space for model building. In $\Gamma_{N}\ll\Gamma_{\phi}$ scenario, there is an RHN matter dominant phase. We calculate the neutrino reheating duration. Utilizing the $N_{\mathrm{RH}}$-$N_{k}$ relation, we arrive at a direct link of the baryon asymmetry with the spectral index, bringing these two important observables on the same plane (figures 2468).
@@
The direct link of $Y_{\mathrm{B}}$ with $n_{s}$ is only possible when we consider neutrino reheating constraints to inflation models. In ref. [1], we draw inflation-compatible lines in BAU parameter space. The inflationary constraints are considered under general assumptions regarding reheating ($N_{k}=60$). In this work, the inflationary observables are calculated via the $N_{\mathrm{RH}}$-$N_{k}$ relation, with $N_{\mathrm{RH}}$ determined during neutrino reheating. In this context, $N_{\mathrm{RH}}$ depends on $T_{\mathrm{RH}}$, leading to a $T_{\mathrm{RH}}$ $n_{s}$ relation, which in turn translates into a $Y_{\mathrm{B}}$ $n_{s}$ relation, considering the analytical limits of non-thermal leptogenesis.
@@
Reheating can break the degeneracy of inflation models. Neutrino reheating is particularly effective in this regard, as it includes predictions for $Y_{\mathrm{B}}$, which are sensitive to $n_{s}$ and thus provide more restrictive constraints. In our considered inflation models, $Y_{\mathrm{B}}$ changes rapidly with $n_{s}$. Bringing $Y_{\mathrm{B}}$ together with $(n_{s},r)$ constrains parameter space to very narrow region near certain $n_{s}$ (marked by the red points in figures 2468). For example, Starobinksy model predicts a line agrees with observation from $n_{s}=0.951$ to $n_{s}=0.964$. In neutrino reheating with $Y_{\mathrm{B}}$ constraint, the viable parameter space is tightly positioned around $n_{s}=0.962$ with other parameters around the values shown in table 1. The same happens for other models. In addition to the Starobinsky model, we find the Coleman-Weinberg inflation with $v=30\mathrm{M}_{\mathrm{P}}$, natural inflation with $f=6\mathrm{M}_{\mathrm{P}}$ consistent with constraints from both inflation and baryon asymmetry. The rest models are excluded by their predictions on $r$.
@@
Connecting inflation with neutrino physics through non-thermal leptogenesis with direct inflaton-RHN coupling offers a simple, natural and testable framework describing the early Univserse. The well-motivated components—inflation, type-I seesaw, and leptogenesis—provide straightforward and elegant solutions to several open questions in cosmology and particle physics. This approach naturally incorporates neutrino reheating, leaving no ambiguity regarding the early history of the Universe. Predictions from neutrino reheating place $Y_{\mathrm{B}}$ and $n_{s}$ on the same plane, providing specific predictions that help break the degeneracy of inflationary models. There remains much to explore in this direction, including the incorporation of neutrino flavor structure and refining the evolution process. We will address these topics in future work and look forward to upcoming cosmological and neutrino data to further test this framework.

### v3

@@ 4.4 Example 4: Natural inflation
$V=\Lambda^{4}\left(1+\cos\frac{\phi}{f}\right)\,.$ (46)
@@
Planck 2018 central value of $n_{s}=0.9649$ gives $f>5.34M_{\mathrm{P}}$. A smaller $n_{s}$ renders a smaller $f$ lower limit. We can express $H_{k}$ as a function of $n_{s}$ with the help of eq. (52) and eq. (54). Similarly, we can get $V_{\mathrm{end}}$ as a function of $n_{s}$ with eq. (50) eq. (51) and $H_{k}$. Note that all these quantities are also functions of $f$. The overall scale $\Lambda$ can be fixed by the scalar perturbation amplitude, which is at $\mathcal{O}(10^{16})$ GeV.
@@
In this limit, $N_{\mathrm{RH}}=0$ and the $N_{\mathrm{RH}}$-$N_{k}$ relation reduces to (10), from which we find a prediction of $f$ as a function of $n_{s}$. It allows us to determine $M_{\phi}$ and also the tensor-to-scalar ratio. $N_{\mathrm{RH}}=0$ also allows us to equate $\rho_{\mathrm{end}}$ with $\rho_{\mathrm{RH}}$. Hence, we find a relation of $T_{\mathrm{RH}}$ with $n_{s}$. The predictions are plotted on the $n_{s}$ $r$ plane in figure 12 as the blue curve. The intersection with the Planck-BK18 $2\sigma$ region allows us to locate the viable parameter space:
$f\in[5.09,8.26]M_{\mathrm{P}}$
$T_{\mathrm{RH}}\in[2.52,2.74]\times 10^{15}\mathrm{GeV}$
$M_{\phi}\in[1.23,1.44]\times 10^{13}\mathrm{GeV}\,.$
The predicted baryon asymmetry are from $1.00\times 10^{-4}$ to $1.09\times 10^{-4}$, several orders larger than observed.
@@
[…]
$M_{N},~{}T_{\mathrm{RH}},~{}f,~{}n_{s}\,.$
@@
Figure 10: Predictions in natural inflation with neutrino reheating in the $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. Other shadings are the same as in the figure 1.
@@
As in previous examples, we first choose to work with $f=6M_{\mathrm{P}},10M_{\mathrm{P}}$ and $M_{N}=10^{11},10^{9}$ GeV. The results are shown in figure 10. We observe that the lines corresponding to $f=10M_{\mathrm{P}}$ overlap with the observed $n_{s}$ $1\sigma$ region, whereas the lines corresponding to $f=6M_{\mathrm{P}}$ do not overlap at all. A larger $f$ corresponds to a lower $T_{\mathrm{RH}}$, a smaller $N_{k}$ and a smaller $\omega_{\mathrm{RH}}$. Different $M_{N}$ lines converge in the $\omega_{\mathrm{RH}}$ $n_{s}$ plane. As the universe approaches radiation domination ($\omega=1/3$), the role of $M_{N}$ becomes increasingly irrelevant. Our results confirm that requiring $\omega_{\mathrm{RH}}\leq 1/3$ gives $r\geq 0.05$ [11], and are compatible with updated results in [53].
Figure 11: The full parameter space in natural inflation with neutrino reheating in the $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. In the left panel, horizontal and vertical dashed lines indicate the upper limits of $T_{\mathrm{RH}}$ and $M_{N}$ respectively, determined by different values of $f$. Other shadings are the same as in the figure 2.
Figure 12: Predictions in natural inflation with neutrino reheating. The red dotted lines show the predicted tensor-to-scalar ratio in the $\Gamma_{N}\ll\Gamma_{\phi}$ scenario, with their solid portions marking the range of spectral indices $n_{s}$ that yield sufficiently large $Y_{\mathrm{B}}$. The blue curve is the prediction in the $\Gamma_{N}\gg\Gamma_{\phi}$ scenario. Other shadings are the same as in figure 3.
The complete parameter space is shown in figures 11 and 12. In the left panel of figure 11, model-dependent constraints appear as dashed lines representing upper bounds on both $T_{\mathrm{RH}}$ and $M_{N}$. The viable parameter space for generating sufficient baryon asymmetry corresponds to the upper-left white triangular region bounded by the red curve and the dashed $T_{\mathrm{RH}}$ constraint. In the right panel of figure 11, each red curve terminates at an upper bound determined by (21). Combining the constraints from figures 11 and 12, we identify the viable range for the parameter $f$ as $[5.09,8.26]~{}M_{\mathrm{P}}$. This range predicts a sufficient baryon asymmetry and agrees with the Planck-BK18 $n_{s}$-$r$ constraint at the $2\sigma$ confidence level. The degeneracy in $f$ can be broken by requiring the correct observed value of $Y_{\mathrm{B}}$, as demonstrated in the right panel of figure 11. While the predicted baryon asymmetry $Y_{\mathrm{B}}$ favors larger values of $n_{s}$, consistent with ACT data, including ACT constraint will exclude all natural inflation with neutrino reheating scenarios at $2\sigma$ confidence level. In the natural inflation model, as $f$ grows, both $n_{s}$ and $r$ grow. The tensor-to-scalar ratio is too large to be compatible with ACT data. Taking into account other effects such as dissipative effects [66, 67] or modified gravity  [68, 69, 70, 71, 72, 73] may suppress $r$ and render a natural inflation model compatible with constraints.
@@ 5 Discussions and conclusions
In this work, we continue investigating the intriguing possibility of connecting inflation and neutrino physics through non-thermal leptogenesis [1]. We assume a direct inflaton-RHN coupling. The RHNs decay reheats the universe through their subsequent decays to SM particles and we called it “neutrino reheating”. Depending on the relative sizes of $\Gamma_{N}$ and $\Gamma_{\phi}$, we consider two scenarios of neutrino reheating. In the $\Gamma_{N}\gg\Gamma_{\phi}$ scenario, the RHNs decay instantaneously, corresponding to $N_{\mathrm{RH}}=0$. We find predictions on $n_{s}$, $T_{\mathrm{RH}}$ and hence $Y_{\mathrm{B}}$ for each model. The predicted $Y_{\mathrm{B}}$ are at $\mathcal{O}(10^{-4})$ to $\mathcal{O}(10^{-5})$, which are much larger than the observed value and leave much space for model building. In the $\Gamma_{N}\ll\Gamma_{\phi}$ scenario, there is an RHN matter dominant phase. We calculate the duration of neutrino reheating. Utilizing the $N_{\mathrm{RH}}$-$N_{k}$ relation, we establish a direct connection between the baryon asymmetry and the spectral index, bringing these two important observables onto the same plane (see right panels of figures 25811).
@@
The direct link of $Y_{\mathrm{B}}$ with $n_{s}$ is only possible when we consider neutrino reheating constraints to inflation models. In ref. [1], we draw inflation-compatible lines in BAU parameter space. Inflationary constraints are considered under general assumptions with respect to reheating ($N_{k}=60$). In this work, the inflationary observables are calculated via the $N_{\mathrm{RH}}$-$N_{k}$ relation, with neutrino reheating determined $N_{\mathrm{RH}}$. In this context, $N_{\mathrm{RH}}$ depends on $T_{\mathrm{RH}}$, leading to a $T_{\mathrm{RH}}$ $n_{s}$ relation, which in turn translates into a $Y_{\mathrm{B}}$ $n_{s}$ relation, considering the analytical limits of non-thermal leptogenesis.
@@
Reheating can break the degeneracy of inflation models. Neutrino reheating is particularly effective in this regard, as it includes predictions for $Y_{\mathrm{B}}$, which are very sensitive to $n_{s}$ and thus provide more restrictive constraints. In our considered inflation models, $Y_{\mathrm{B}}$ changes rapidly with $n_{s}$. Bringing $Y_{\mathrm{B}}$ together with $(n_{s},r)$ constrains the parameter space greatly as shown in figures 36912. Our results exhibit partial agreement with Planck-BK18 data for the parameter range $v\sim[13,47]M_{\mathrm{P}}$ in the Coleman-Weinberg inflation, and $f\in[5.09,8.26]~{}M_{\mathrm{P}}$ in the natural inflation model.
The reheating temperature $T_{\mathrm{RH}}$ is related to the spectral tilt through the $N_{\mathrm{RH}}$-$N_{k}$ relation with $N_{\mathrm{RH}}$ determined by neutrino reheating. It is highly sensitive to $n_{s}$ and spans several orders of magnitude given the $n_{s}$ constraint. It also leads to rapid changes of the baryon asymmetry in the limit $\Gamma_{N}\ll\Gamma_{\phi}$. With the precise determination of $Y_{\mathrm{B}}$, $T_{\mathrm{RH}}$ is found to range from $10^{4}$ to $10^{9}$ GeV in the models considered. In the limit $\Gamma_{N}\gg\Gamma_{\phi}$, $T_{\mathrm{RH}}$ is much higher, around $10^{15}$ GeV, which may lead to a potential gravitino problem in supersymmetric models [74, 75, 76]. Possible solutions, other than lower $T_{\mathrm{RH}}$, can be found in [77] and references therein.
@@
Connecting inflation with neutrino physics through non-thermal leptogenesis with direct inflaton-RHN coupling offers a simple, natural, and testable framework describing the early univserse. The well-motivated components—inflation, type-I seesaw, and leptogenesis—provide straightforward and elegant solutions to several open questions in cosmology and particle physics. This approach naturally incorporates neutrino reheating, leaving no ambiguity regarding the early history of the universe. The predictions of neutrino reheating place $Y_{\mathrm{B}}$ and $n_{s}$ on the same plane, providing specific predictions that help break the degeneracy of inflationary models. There remains much to explore in this direction, including the incorporation of neutrino flavor structure and the refinement of the evolution process. We will address these topics in future work and look forward to upcoming cosmological and neutrino data to further test this framework.